You have decided to advance refund $10,000,000 of outstanding debt that is callable in five years. The interest rate on these bonds is 8 percent. You can issue new bonds at 6 percent. For every dollar of new debt issued, you will incur a 5 percent issuance cost. Interest payments on the present issue are $800,000 per year with no scheduled principal payments. How much new debt needs to be issued to realize the defeasance of the present issue?
Answer #1

Answol compute the present value of the collable debt (int) : year-o pousent value = 800,000 X (1+0.06)? = 800,000 X 0.9433 =present blue = 800,000 X (1+0.06) -800,000 X 0.147 = 597,600 => 154,640 + 11,992 + 671,688 +633,670 4597,600 = 3,369,592 Pres